OpenStudy (anonymous):

s=1/2gt^2 solve for g and please show me how you got it

OpenStudy (anonymous):

\[s=\frac{1}{2gt^2}\]? like that?

OpenStudy (anonymous):

or like \[s=\frac{1}{2}gt^2\]

OpenStudy (anonymous):

sorry the second one

OpenStudy (anonymous):

ok then two steps: 1) multiply by 2 2) divide by \[t^2\]

OpenStudy (anonymous):

\[s=\frac{1}{2}gt^2\] \[2s=gt^2\] \[\frac{2s}{t^2}=g\]

OpenStudy (anonymous):

okay, thanks!!